# Multiclass Local Calibration With the Jensen-Shannon Distance

### v1

@@ C.2 Local Calibration and Proximity Bias
$\left\|\text{freq}(S_{1})-\text{freq}(S_{2})\right\|_{1}\leq 2\varepsilon+\sqrt{\frac{2\log(4C/\delta)}{\min(|S_{1}|,|S_{2}|)}}+\left\|\text{conf}(S_{1})-\text{conf}(S_{2})\right\|_{1}$
Then, under the local calibration assumption $\|f(\mathbf{x}_{j})-\hat{\theta}(\mathbf{y}_{j}\mid\mathbf{x}_{j})\|_{1}\leq\varepsilon$ for all $j\in D$, we have:
$\lim_{n\to\infty}\Pr\!\Big(\|\mathrm{freq}(S_{1})-\mathrm{freq}(S_{2})\|_{1}\leq 2\varepsilon\Big)=1.$ (28)
Proof.
@@
Then, by the Lebesgue differentiation theorem (see e.g. Theorem 1.3 in [22]), for sufficiently small $r_{n}$ and $w_{i}$,
@@
Since $\mathbb{E}[|S_{s}|]\to\infty$ as $n\to\infty$, with probability tending to one,
$|S_{s}|\geq(1-\eta)\,\mathbb{E}[|S_{s}|]\to\infty,\qquad\text{i.e.,}\qquad\Pr(|S_{s}|\to\infty)\to 1.$ (30)
@@
$\lim_{n\to\infty}\Pr\!\Big(\|\mathrm{freq}(S_{1})-\mathrm{freq}(S_{2})\|_{1}\leq 2\varepsilon\Big)=1.$ (31)
@@
The argument readily extends to the case where each proximity-based sub-bin $S_{s}$ is a finite union of disjoint regions $\{R_{s,k}\}_{k=1}^{K_{s}}$. Under the same regularity and consistency assumptions applied component-wise (shrinking diameters and diverging per-component sample sizes), the concentration and continuity arguments hold uniformly over components, and the aggregate deviation remains bounded by $2\varepsilon$ with maximum probability. We restrict the proof to a single region per sub-bin for notational simplicity.
@@ C.3 Illustrative Example
This recalibrated probability equals a weighted average of the empirical frequencies of regions $b$ and $c$. As such, it cannot simultaneously correct both, and whichever side is under/overconfident remains miscalibrated after recalibration. The magnitude of miscalibration aggravates when one region is underconfident and the other is overconfident. Such heterogeneous calibration errors do not exclusively occur when regions with similar predicted confidences differ in density but can also arise in presence of differences in class overlap or representation smoothness, as commonly observed in deep neural networks [5, 13]. Consequently, grouping them within the same density–confidence bin averages incompatible local behaviors, masking or amplifying miscalibration. While narrower bins could mitigate this effect, they quickly become sample-inefficient and computationally demanding.
@@ D.1 Training of Classifiers
We use a ResNet-50 architecture initialized with IMAGENET1K_V2 pre-trained weights. A dropout layer with rate $0.2$ is appended to the final backbone layer, followed by a linear classification head. During fine-tuning, all layers are frozen except for the last backbone block and the classification head. Optimization is performed for $9$ epochs using the Adam optimizer [10] with a learning rate of $3\!\times\!10^{-4}$.
@@ D.2 Training of LoCal Nets
We report here all training configurations for LCN in the calibration experiments.
@@
The LCN operates in a residual fashion. Given the intermediate representations $\phi(\mathbf{x})$ extracted from a pre-trained backbone, let $\phi_{\mathrm{PCA}}(\mathbf{x})$ denote the reduced feature representation obtained via Principal Component Analysis (PCA). The LCN processes $\phi(\mathbf{x})$ through its hidden layer to produce refined features $\tilde{\phi}_{\mathrm{PCA}}(\mathbf{x})$ and logits $\tilde{g}(\mathbf{x})$. The final representations $\phi_{\mathrm{PCA}}^{\prime}(\mathbf{x})$ and $g^{\prime}(\mathbf{x})$ are obtained through a weighted residual combination:
@@
where $w_{\phi},b_{\phi},w_{g}$ and $b_{g}$ are learnable scalar weights and biases that adaptively control the contribution of the original features and logits, respectively. Weights are initialized as $1$ and biases are randomly sampled from normal distributions with $0.$ location and $0.01$ scale parameters. The residual formulation provides strong initialization for LCN outputs, enables preserving the semantic content of the backbone features while introducing locally calibrated corrections, and improves both stability and convergence to meaningful solutions.
@@
The LCN is implemented as a fully connected network with a single hidden layer of size $64$ and dropout rate $0.3$. It has two output heads: one of dimension $10$ (corresponding to the number of classes) and one of dimension $50$, used for the PCA-reduced feature representations. The loss weighting hyperparameter $\lambda$ is set to $1$, ensuring equal contribution of both components of the objective. The fixed kernel bandwidth $\gamma$ is set to $10$, chosen to be as small as possible to preserve locality while maintaining stable convergence of the cross-entropy component of the loss, as excessively small values lead to training collapse. This choice is validated empirically using a held-out validation set. Optimization is performed using the Adam optimizer with learning rate $1\!\times\!10^{-3}$, for $22$ epochs and a batch size of $1024$.
@@
The LCN uses a fully connected architecture with one hidden layer of size $128$ and dropout rate $0.5$. As before, it has two output heads: one of dimension $100$ (matching the number of classes) and one of dimension $50$ for the PCA-reduced features. We set $\lambda=1$ for equal loss weighting and fix $\gamma=10$ for consistency with the other datasets. In this case, slightly smaller bandwidth values were found feasible, but $\gamma=10$ was retained for coherence across experiments. Optimization uses Adam with learning rate $1\!\times\!10^{-3}$, trained for $30$ epochs with a batch size of $1024$.
@@
The LCN is a single-hidden-layer fully connected network with hidden dimension $256$ and dropout rate $0.3$. It includes two output heads: one of dimension $8$ (the number of classes) and one of dimension $50$ for the PCA-reduced representations. The hyperparameter $\lambda$ is set to $1$, and the bandwidth $\gamma$ is fixed at $10$, following the same locality–stability trade-off principle described above, validated via a held-out set. Optimization uses Adam with a learning rate of $1\!\times\!10^{-3}$ for $60$ epochs with a batch size of $1024$.
@@ D.3 Metrics
We employ two kernel-based calibration metrics: the multiclass Local Calibration Error (LCE) and its maximum variant (MLCE). To extend $LCE$ to the multiclass setting, we adopt a class-wise formulation analogous to that used for binning-based metrics. Specifically, $f_{c}(\mathbf{x})$ is partitioned into $15$ confidence-based bins, and for each fixed class, we use the corresponding bins to identify the neighborhood of each anchor point for kernel estimation. Bins with fewer than 20 elements were discarded to prevent using unstable kernel estimates. For each $i\in D$, the $LCE$ is computed as the absolute difference between kernel-weighted estimates of predicted confidences and empirical labels of instances in the same confidence bin. Per-sample deviations are then averaged, and the values for each class are combined using priors to obtain the final $LCE$. The kernel bandwidth parameter was set to $\gamma=10$, consistent with the bandwidth used during the training of the LCN.

### v2

@@ C.2 Local Calibration and Proximity Bias
$\left\|\text{freq}(S_{1})-\text{freq}(S_{2})\right\|_{1}\leq 2\varepsilon+\sqrt{\frac{2\log(\frac{4C}{\delta})}{\min(|S_{1}|,|S_{2}|)}}+\left\|\text{conf}(S_{1})-\text{conf}(S_{2})\right\|_{1}$
Then, under the local calibration assumption $\|f(\mathbf{x}_{j})-\hat{\theta}(\mathbf{y}_{j}\mid\mathbf{x}_{j})\|_{1}\leq\varepsilon$ for all $j\in D$, we will prove that:
$\forall\;\xi>0:\lim_{n\to\infty}\Pr\Big(\|\mathrm{freq}(S_{1})-\mathrm{freq}(S_{2})\|_{1}\leq 2\varepsilon+\xi\Big)=1$ (47)
@@
Then, by the Lebesgue differentiation theorem (see e.g. Theorem 1.3 in (Stein and Shakarchi, 2009)), for sufficiently small $r_{n}$ and $w_{i}$,
@@
Since $\mathbb{E}[|S_{s}|]\to\infty$ as $n\to\infty$, with probability tending to one:
$|S_{s}|\geq(1-\eta)\,\mathbb{E}[|S_{s}|]\to\infty$ (49)
@@
$\forall\;\xi>0:\lim_{n\to\infty}\Pr\Big(\|\mathrm{freq}(S_{1})-\mathrm{freq}(S_{2})\|_{1}\leq 2\varepsilon+\xi\Big)=1$ (50)
@@
The argument readily extends to the case where each proximity-based sub-bin $S_{s}$ is a finite union of disjoint regions $\{R_{s,k}\}_{k=1}^{K_{s}}$. Under the same regularity and consistency assumptions applied component-wise (shrinking diameters and diverging per-component sample sizes), the concentration and continuity arguments hold uniformly over components, and the aggregate deviation remains bounded by $2\varepsilon$ up to vanishing terms with maximum probability. We restrict the proof to a single region per sub-bin for notational simplicity.
@@ C.3 Illustrative Example
This recalibrated probability equals a weighted average of the empirical frequencies of regions $b$ and $c$. As such, it cannot simultaneously correct both, and whichever side is under/overconfident remains miscalibrated after recalibration. The magnitude of miscalibration aggravates when one region is underconfident and the other is overconfident. Such heterogeneous calibration errors do not exclusively occur when regions with similar predicted confidences differ in density but can also arise in presence of differences in class overlap or representation smoothness, as commonly observed in deep neural networks (Guo et al., 2017; Kull et al., 2019). Consequently, grouping them within the same density–confidence bin averages incompatible local behaviors, masking or amplifying miscalibration. While narrower bins could mitigate this effect, they quickly become sample-inefficient and computationally demanding.
@@ D.1 Training of Classifiers
We use a ResNet-50 architecture initialized with IMAGENET1K_V2 pre-trained weights. A dropout layer with rate $0.2$ is appended to the final backbone layer, followed by a linear classification head. During fine-tuning, all layers are frozen except for the last backbone block and the classification head. Optimization is performed for $9$ epochs using the Adam optimizer (Kingma and Ba, 2015) with a learning rate of $3\!\times\!10^{-4}$.
@@ D.2 Training of LoCal Nets
We report here all training configurations for LN in the calibration experiments.
@@
The LN operates in a residual fashion. Given the intermediate representations $\phi(\mathbf{x})$ extracted from a pre-trained backbone, let $\phi_{\mathrm{PCA}}(\mathbf{x})$ denote the reduced feature representation obtained via Principal Component Analysis (PCA). The LN processes $\phi(\mathbf{x})$ through its hidden layer to produce refined features $\tilde{\phi}_{\mathrm{PCA}}(\mathbf{x})$ and logits $\tilde{g}(\mathbf{x})$. The final representations $\phi_{\mathrm{PCA}}^{\prime}(\mathbf{x})$ and $g^{\prime}(\mathbf{x})$ are obtained through a weighted residual combination:
@@
where $w_{\phi},b_{\phi},w_{g}$ and $b_{g}$ are learnable scalar weights and biases that adaptively control the contribution of the original features and logits, respectively. Weights are initialized as $1$ and biases are randomly sampled from normal distributions with $0.$ location and $0.01$ scale parameters. The residual formulation provides strong initialization for LN outputs, enables preserving the semantic content of the backbone features while introducing locally calibrated corrections, and improves both stability and convergence to meaningful solutions.
@@
The LN is implemented as a fully connected network with a single hidden layer of size $64$ and dropout rate $0.3$. It has two output heads: one of dimension $10$ (corresponding to the number of classes) and one of dimension $50$, used for the PCA-reduced feature representations. The loss weighting hyperparameter $\lambda$ is set to $1$, ensuring equal contribution of both components of the objective. The fixed kernel bandwidth $\gamma$ is set to $10$, chosen to be as small as possible to preserve locality while maintaining stable convergence of the cross-entropy component of the loss, as excessively small values lead to training collapse. This choice is validated empirically using a held-out validation set. Optimization is performed using the Adam optimizer with learning rate $1\!\times\!10^{-3}$, for $22$ epochs and a batch size of $1024$.
@@
The LN uses a fully connected architecture with one hidden layer of size $128$ and dropout rate $0.5$. As before, it has two output heads: one of dimension $100$ (matching the number of classes) and one of dimension $50$ for the PCA-reduced features. We set $\lambda=1$ for equal loss weighting and fix $\gamma=10$ for consistency with the other datasets. In this case, slightly smaller bandwidth values were found feasible, but $\gamma=10$ was retained for coherence across experiments. Optimization uses Adam with learning rate $1\!\times\!10^{-3}$, trained for $30$ epochs with a batch size of $1024$.
@@
The LN is a single-hidden-layer fully connected network with hidden dimension $256$ and dropout rate $0.3$. It includes two output heads: one of dimension $8$ (the number of classes) and one of dimension $50$ for the PCA-reduced representations. The hyperparameter $\lambda$ is set to $1$, and the bandwidth $\gamma$ is fixed at $10$, following the same locality–stability trade-off principle described above, validated via a held-out set. Optimization uses Adam with a learning rate of $1\!\times\!10^{-3}$ for $60$ epochs with a batch size of $1024$.
@@ D.3 Metrics
We employ two kernel-based calibration metrics: the multiclass Local Calibration Error (LCE) and its maximum variant (MLCE). To extend $LCE$ to the multiclass setting, we adopt a class-wise formulation analogous to that used for binning-based metrics. Specifically, $f_{c}(\mathbf{x})$ is partitioned into $15$ confidence-based bins, and for each fixed class, we use the corresponding bins to identify the neighborhood of each anchor point for kernel estimation. Bins with fewer than 20 elements were discarded to prevent using unstable kernel estimates. For each $i\in D$, the $LCE$ is computed as the absolute difference between kernel-weighted estimates of predicted confidences and empirical labels of instances in the same confidence bin. Per-sample deviations are then averaged, and the values for each class are combined using priors to obtain the final $LCE$. The kernel bandwidth parameter was set to $\gamma=10$, consistent with the bandwidth used during the training of the LN.
@@ D.4 Hardware and Training Time
Appendix E Detailed Results
Dataset Method $LCE\,\downarrow$ $MLCE\,\downarrow$ $ECCE\,\downarrow$ $ECE\,\downarrow$ $ACC\,\uparrow$ $NLL\,\downarrow$
cifar10 LN $\mathbf{.0079\pm.0003}$ $\mathbf{.6279\pm.0157}$ $017\pm.0001 .00$$048\pm.0003$ $888\pm.002 .$$347\pm.002$
DC $0099\pm.0003 .$$8133\pm.0195 .$$\mathbf{.0011\pm.0001}$ $\mathbf{.0037\pm.0001}$ $890\pm.002 .$$\mathbf{.332\pm.007}$
KO $\underline{.0086\pm.0003}$ $8190\pm.0183 .$$\underline{.0012\pm.0001}$ $\underline{.0045\pm.0002}$ $\underline{.893\pm.001}$ $340\pm.002$
KC $0095\pm.0005 .$$\underline{.7923\pm.0269}$ $015\pm.0003 .00$$054\pm.0004$ $\mathbf{.893\pm.001}$ $\underline{.340\pm.003}$
PS $0180\pm.0005 .$$8542\pm.0224 .$$014\pm.0001 .00$$085\pm.0003$ $884\pm.001 .$$466\pm.004$
IR $0115\pm.0003 .$$7978\pm.0162 .$$014\pm.0002 .00$$047\pm.0001$ $884\pm.001 .$$364\pm.008$
TS $0127\pm.0006 .$$8200\pm.0191 .$$036\pm.0008 .00$$072\pm.0010$ $884\pm.001 .$$362\pm.008$
NC $0154\pm.0005 .$$8738\pm.0238 .$$0065\pm.0004 .$$0153\pm.0006 .$$884\pm.001 .$$494\pm.018$
[…]
DC $\underline{.0027\pm.0001}$ $8117\pm.0121 .$$\mathbf{.0007\pm.0001}$ $\underline{.0019\pm.0001}$ $\mathbf{.690\pm.002}$ $\underline{1.154\pm.009}$
KO $0043\pm.0001 .$$7575\pm.0101 .$$013\pm.0001 .00$$044\pm.0001$ $679\pm.003$ $1.352\pm.009$
KC $0043\pm.0001 .$$7594\pm.0055 .$$013\pm.0001 .00$$044\pm.0002$ $679\pm.004$ $1.351\pm.007$
PS $0053\pm.0001 .$$\underline{.7031\pm.0409}$ $011\pm.0001 .00$$033\pm.0001$ $670\pm.002$ $1.618\pm.007$
IR $0031\pm.0001 .$$8049\pm.0116 .$$\underline{.0007\pm.0001}$ $0020\pm.0001 .$$670\pm.002$ $1.437\pm.029$
TS $0034\pm.0001 .$$8239\pm.0173 .$$013\pm.0001 .00$$024\pm.0001$ $670\pm.002$ $1.277\pm.009$
NC $0032\pm.0001 .$$8250\pm.0176 .$$017\pm.0001 .00$$039\pm.0002$ $670\pm.002$ $1.502\pm.036$
tissue LN $\mathbf{.0144\pm.0012}$ $7293\pm.0359 .$$0050\pm.0011 .$$0100\pm.0015 .$$630\pm.001$ $\mathbf{1.012\pm.003}$
DC $0276\pm.0013 .$$9638\pm.0104 .$$\underline{.0021\pm.0005}$ $\mathbf{.0062\pm.0007}$ $617\pm.003$ $1.052\pm.009$
KO $0177\pm.0006 .$$\mathbf{.6913\pm.0349}$ $0047\pm.0006 .$$0144\pm.0006 .$$\underline{.632\pm.001}$ $1.027\pm.002$
KC $\underline{.0165\pm.0008}$ $\underline{.7179\pm.0250}$ $0043\pm.0006 .$$0126\pm.0010 .$$\mathbf{.632\pm.001}$ $\underline{1.021\pm.002}$
PS $0418\pm.0018 .$$9521\pm.0043 .$$\mathbf{.0013\pm.0001}$ $0135\pm.0012 .$$603\pm.008$ $1.180\pm.008$
[…]
TS $0413\pm.0034 .$$9620\pm.0107 .$$0127\pm.0025 .$$0196\pm.0036 .$$603\pm.008$ $1.112\pm.023$
NC $0768\pm.0019 .$$9695\pm.0127 .$$0308\pm.0016 .$$0725\pm.0020 .$$603\pm.008$ $2.100\pm.101$
Table 3: Results for all experiments over five seeds.
Table 3 reports all the results for Q1, Q2 and Q3.